# Bright single-photon source in a silicon chip by nanoscale positioning of a color center in a microcavity

### v1

@@ Abstract
We present an all-silicon source of near-infrared linearly-polarized single photons, fabricated by nanoscale positioning of a color center in a silicon-on-insulator microcavity. The color center consists of a single W center, created at a well-defined position by Si+ ion implantation through a 150 nm-diameter nanohole in a mask. A circular Bragg grating cavity resonant with the W’s zero-phonon line at 1217 nm is fabricated at the same location as the nanohole. Under above-gap continuous-wave excitation, a very clean photon antibunching behavior () is observed over the entire power range, which highlights the absence of parasitic emitters. Purcell-enhancement of W’s zero-phonon emission provides both a record-high photoluminescence count rate among Si color centers (ca counts/s) and apparent Debye-Waller factor around 99%. We also demonstrate the triggered emission of single photons with 93% purity under weak pulsed laser excitation. At high pulsed laser power, we reveal a detrimental effect of repumping processes, that could be mitigated using selective pumping schemes in the future. These results represent a major step towards on-demand sources of indistinguishable near-infrared single photons within silicon photonics chips.
keywords:
[…]
@@ Introduction
The generation, processing, and detection of single photons have enabled pioneering investigations in quantum physics, paving the way to groundbreaking technologies. Silicon quantum photonics, namely the processing of photonic quantum states in a silicon chip, is at the forefront of this revolution.1, 2 However, the potential of silicon photonics for large-scale quantum applications is limited by the difficulty to deterministically generate arbitrary large entangled photon states on-chip. A key requirement to overcome this limitation is to embed sources of indistinguishable single photons operating on demand. While integrated sources of entangled photon pairs based on spontaneous four-wave mixing have enabled major demonstrations,2 these sources suffer from an inherent interplay between efficiency and multi-photon pairs emission. Alternative approaches rely on fiber-coupling of an external source,3, 4 and hybrid integration of quantum emitters,5, 6, 7 albeit with limited efficiency due to coupling losses.
@@
In this context, near-infrared color centers in silicon have emerged as a promising path towards deterministic single-photon sources (SPS) circumventing coupling losses. Due to their emission wavelength in the telecom bands, these point defects also appear as potential sources for chip-based quantum communications. Since the observation of photon antibunching for single color centers in 2020,8, 9 tremendous efforts have been devoted to optimizing their elaboration processes,10, 11, 12, 13, 14 and exploring their photophysical properties.15, 16, 17, 18, 19 In order to build a SPS, it is necessary to embed the emitter in a photonic cavity to achieve on-demand emission of single photons in a well-defined optical mode using the Purcell effect.20 Several single-mode sources in silicon have been demonstrated, based on a T center,21, 22 G center,23 G⋆ center,24 and erbium dopant.25 However, in spite of remarkable improvement of emission properties owing to a high Purcell factor, these sources suffer from significant limitations. The T center and erbium dopant are inherently slow emitters,21, 22 the G center has a poor quantum efficiency,26, 27, 23 and the fabrication process of the G⋆ center is not selective,15 resulting in poor purity due to the presence of parasitic defects.24 Moreover, in these first experiments the position of the emitter in the cavity is not controlled, hindering both the optimal exploitation of the Purcell effect and the reproducible fabrication of the SPS.
@@
In recent work, Purcell enhancement has been demonstrated for an ensemble of W centers in circular Bragg grating (CBG) cavities in a silicon-on-insulator (SOI) film.28, 29 On the one hand, these results have positioned the W center as a promising candidate to build a SPS, owing to its high quantum efficiency () and relatively fast zero-phonon emission ( ns) in bulk silicon. 28 On the other hand, it confirms that CBG cavities, which have been widely used in other material systems,30, 31, 32, 33, 34, 35 are also attractive for building SPS on SOI chips.
@@
In this letter, we present a SPS fabricated by nanoscale positioning of a single W center in a SOI CBG cavity. An array of W centers is created by Si+ ion implantation through a mask patterned with nanoholes,11 then an array of CBG cavities centered at the same location as the nanoholes and resonant with the W zero-phonon line (ZPL) is fabricated. The optical properties of a cavity-coupled W center are studied by photoluminescence (PL) spectroscopy and photon correlation measurements under above-gap continuous-wave (CW) and pulsed excitation. The purity of the source is analyzed as a function of the laser excitation power.
@@
A 480 nm-thick mask of polymethyl methacrylate (PMMA) is deposited onto a commercial piece of SOI wafer. Arrays of nanoholes defined by electron beam lithography are etched into the mask The sample is then broad-beam implanted with Si+ ions (Figure 1(a)). A similar approach has already proven efficient for the fabrication of NV centers at preset positions in diamond,36, 37 and, more recently, G and W centers in silicon.11 The PMMA mask is then removed by stripping with acetone and the sample is annealed at 200 °C for 30 minutes in N2 atmosphere (Figure 1(b)). This second step leads to the activation of the W center,38, 39 whose microscopic structure (consisting of three interstitial Si atoms) is shown in the inset. A characterization of the emitters in unpatterned SOI is presented in Section 3 of the Supporting Information. CBG cavities centered at the same location as the nanoholes are fabricated by electron beam lithography and reactive ion etching (Figure 1(c)). This method is well-tried for the high-yield fabrication of SPS based on III-V semiconductor quantum dots (QD).40, 41 A map of the electromagnetic field in the central disk of the cavity is shown in the inset. After fabrication, the sample is first inspected by scanning electron microscopy, to confirm successful fabrication according to design parameters (see Figure S11 in the Supporting Information).
@@
A sectional view of a CBG cavity containing a single W center is schematically represented in Figure 1(d), with the relevant dimensions. The red double-arrow represents the orientation of the transition dipole along the direction.16 It is well-known for CBG cavities that the Purcell enhancement factor , defined as the ratio between the emission rate in the cavity and bulk semiconductor, is heavily dependent on the position of the emitter.40 This sensivity was investigated for our emitter-cavity system using numerical finite-difference time-domain simulations (FDTD, see Section 4 in the Supporting Information). Figure 1(e) shows the Purcell enhancement factor for the ZPL as a function of the distance between the center of the cavity and the position of the dipole, in the coordinate system defined in Figure 1(d). A maximum enhancement factor of 13 is obtained at the center of the cavity. decreases rapidly against and reaches 0.3 for nm, emphasing the requirement for high positioning accuracy. Considering the simulated lateral spread of the ions implanted through a nanohole of 150 nm diameter, we estimate that the W centers have near-unity probability to be located less than 120 nm away from the center of the cavity, i.e. in the vicinity of the main antinode of the electromagnetic field (see Section 2 in the Supporting Information).
@@
The collection efficiency for a cavity-coupled W center was also investigated using FDTD, considering the configuration of our measurement setup. The fraction of dipole emission radiated through a numerical aperture of 0.65 at normal incidence is plotted against in Figure 1(e). A collection efficiency of 51 is obtained at the center of the cavity. Note that this value could be increased to nearly with the addition of a high-reflectivity mirror below the cavity.42 The curves also reveal the robustness of the collection efficiency against spatial detuning (with for nm for each three coordinates), except for the "pit" around = 120 nm where the emitter is at an antinode of the electromagnetic field. This feature makes CBG cavities an attractive microstructure to easily increase the photon detection rate from quantum emitters in SOI. Similarly, the collection efficiency is also found robust against spectral detuning (see Figure S7 in the Supporting Information).
@@
After fabrication, the sample is characterized by mapping the PL intensity in the emission range of W centers. On PL maps, CBG cavities that contain W centers appear as bright spots on a black background (Figure 2(a)). As the number of W centers formed through localized ion implantation is a random variable, we have thus to identify the CBG cavities that contain a single W center. Since the PL emission of single W centers, collected under normal incidence, is linearly polarized along either [110] or [1] crystal axes of Si, we identify very easily the CBG cavities that are the most likely to contain a single W emitter, by comparing two polarized-resolved PL maps, obtained for those crossed polarizations (see Section 5 in the Supporting Information). Photon-autocorrelation experiments in the Hanbury Brown and Twiss configuration are then performed to confirm the presence of one and only one W center in the cavity.
We display in Figure 2 a typical set of results obtained for a CBG cavity containing a single W center under above-gap CW excitation (see Section 1 in the Supporting Information for details about the measurement configuration). A measurement of the polarization diagram of the PL evidences linear polarization along the [110] axis of the silicon crystal with 97.2 0.7 visibility (see Figure S9 in the Supporting Information). We observe of a clean antibunching dip in the photon-autocorrelation histogram shown in Figure 2(c, d). It is worth emphasizing that this low value of is obtained without any correction for detectors dark counts or background emission. A bunching behaviour () is observed at non-zero delay. This phenomenon, already reported for single W centers,16, 11 is attributed to a metastable state coupled to the excited state of the emitter.
In order to highlight cavity effects on the single W center emission, we also show in Figure 2(b) a study of the PL intensity as a function of the CW pump power. We observe a well defined saturation behaviour and a count rate at saturation around Mcps. This value is remarkably high compared to previously published results for other color centers in SOI, as shown in Table 1. Importantly, we observe a very low over the full range of pump powers up to the saturation plateau ( is equal to 0.03 at 3 and to 0.06 at 9 ), unlike most previous reports. This important observation confirms that our fabrication process does not produce parasitic emitters, besides the desired isolated W center. The beneficial effect of the CBG cavity is most obviously evidenced by comparing the latter PL count rate at saturation to the one of isolated W centers in unpatterned SOI which is typically around 3.2 0.8 kcps after correction for background counts rate (see Section 3 in the Supporting Information). This represents an enhancement factor of around 400 for the PL count rate thanks to cavity effects. As shown below, this increase results from the combination of Purcell-enhancement of zero-photon emission into the CBG cavity mode and the small divergence of the radiation pattern of this mode in the far-field, providing efficient collection of zero-phonon emission.
| Em. | Cav. | (kcps) | (ns) | Pos. | Reference | |
| G⋆ | 2D PhC | 20 | 6.7 | 0.30 | No | Redjem et al.24 |
| G | 2D PhC | 20 | 6 | 0.03 | No | Saggio et al. see SI in ref.[23] |
| T | 1D PhC | 20 | 169 | 0.07 | No | Islam et al.21 |
@@
To understand quantitatively cavity effects, we investigated the properties of the cavity mode and the spectral resonance of the emitter-cavity system. The modal properties have been probed by performing reflection spectroscopy under white light illumination. As shown in Figure 3(b), the reflectivity curve features a dip around 1218 nm, fitting to a Lorentzian profile with a full width at half maximum (FWHM) nm, corresponding to a quality factor . This value is in good agreement with the quality factor estimate of 167 obtained from FDTD simulations. The CBG cavity mode is therefore nearly perfectly matched to the ZPL emission of zero-phonon emission line of W centers in unstrained silicon,16 as confirmed from a comparison of reflectivity and PL emission spectra in Figure 3(a). Noticeably, this PL spectrum over the 1140-1450 nm spectral range reveals that zero-phonon emission accounts for of the PL intensity, value to be compared with 39 for isolated W in unpatterned SOI, which unveils a selective enhancement of ZPL emission thanks to the Purcell effect.
@@
We study now the recombination dynamics of the W center to estimate the magnitude of the Purcell effect. The PL decay of the emitter under pulsed excitation at 17.5 W with a 5 MHz repetition rate is displayed in Figure 3(c). The intensity decays nearly mono-exponentially with a characteristic time ns. This value should be compared with the characteristic time of around 34 ns obtained for ensembles of W centers in unpatterned SOI.39, 28, 29 The acceleration of the decay compared to this reference value has two possible origins: Purcell enhancement of zero-phonon emission, and/or activation of extra non-radiative recombination channels due to the interaction between the emitter and its environment. This second possibility cannot be neglected a priori as a significant lifetime dispersion has been reported for single W centers in 60 nm-thick SOI,16 although this dispersion can likely be attributed to surface effects in such a thin SOI layer. A detailed analysis of the photon budget, from emission by W centers to detection by our single photon detectors, provides a lower-bound estimate around 1/18 ns-1 for the zero-phonon emission rate (see Section 6 in the Supporting Information). By comparison with a previous estimate of 1/134 ns-1 in bulk Si,28 we conclude that the zero-phonon emission rate is increased by a factor . Therefore, the acceleration of the PL decay results predominantly (and possibly entirely) from the Purcell effect. Additionally, according to Figure 1(e), we can also conclude that the emitter is positionned less than 100 nm away from the cavity center.
@@
The bright emission of antibunched photons under CW excitation unveils the strong potential of W centers for building a deterministic SPS in SOI. Along this line, we have explored the triggered emission of single photons under pulsed excitation. Figure 4(a) shows uncorrected autocorrelation histograms obtained for different pump powers for a 5MHz repetition rate. In this figure, histograms are normalized such that the area of all peaks would be equal to 1 for a coherent source. The normalized area as a function of time delay , where is the laser repetition period, of each peak is given by the coordinate of the squares on the right axis. In the low power regime, the peak at zero-delay has an area , demonstrating the triggered emission of single-photon pulses with 93 purity. As for the measurement under CW excitation, a bunching behaviour is observed (>1), indicating photoblinking of the source. Note that this effect is a rather common issue for solid-state quantum emitters and is expected for W, as this color center possesses a metastable excited state.16 The area of the peaks follows an exponential law given by:
@@
where and are the average times of stay of the emitter in the "bright" and "dark" states, respectively.43 For = 0.5, 4.4, and 17.5 W, = 2290, 341, and 153 ns, and = 13200, 2200, and 741 ns, pointing towards an acceleration of the blinking with increasing power. Interestingly, the peaks for have a smaller area than expected from the model in the low-power regime. This effect, which has been reported for single InAs QD under above-gap excitation,44 may be related to a two-step transition from the ground state to the excited state, which is indeed predicted by ab-initio calculation.16
@@
The area of the central peak increases with power, as shown in Figure 4(b). For = 17.5 W, the area of the peak reaches 0.51, meaning that about two photons are emitted on average per excitation pulse. A zoom on the peak with 20 ps time bins (Figure 4(c)) reveals an antibunching dip with a characteristic time of 60 ps (that disappears in Figure 4(b) due to rebinning). This antibunching behaviour indicates that the two photons are emitted successively. An explanation is that, under strong above-gap pumping condition, a second exciton can be captured after the emission of a first photon, as reported for other quantum emitters.45 This phenomenon is not surprizing here, due to the optical injection of a large population of carriers by the pump pulse in the central disk of the cavity. As for QD-SPS, we expect to overcome this limitation through the implementation of resonant or quasi-resonant pumping schemes.
@@
Under pulsed excitation at 5MHz frequency, the PL count rate reaches 52 kcps at saturation. This value corresponds, after taking into account transmission losses, to a source efficiency (as measured at the first lens of our collection optics) comprised between 0.03 and 0.06 photon per pulse (see Section 6 of the SI). The main factors limiting at this stage are easily identified: (1) W is in its bright state one sixth of the time, which reduces its average efficiency by the same amount; (2) our CBG cavity funnels only one-half of the emission towards our collection optics; (3) parasitic recombination channels (non-radiative ones and to a lesser extent phonon-assisted recombination) reduce the quantum efficiency of the zero-phonon emission to a value comprised between 0.36 and 0.68.
While point (1) and the issue of blinking has already been discussed, we see that points (2) and (3) can be radically improved by replacing our low-Fp CBG cavity by a high-Fp waveguide-coupled photonic crystal cavity, such as a nanobeam cavity. 46, 21, 22 Proper engineering of light extraction towards the waveguide mode and a large but realistic Purcell factor around 150 would bring the zero-phonon emission rate above 109 s-1. This would ensure that the quantum efficiency of the ZPL, the spontaneous emission factor and the efficiency of the waveguide-coupled SPS are all near unity. Single photon coherence and indistinguishability are also required properties for applications to quantum simulations and computing. Until now, Hong-Ou-Mandel interference have only been demonstrated over relatively narrow time-windows, shorter than the emitter lifetime. 47 Combining such a large Purcell acceleration of zero-phonon emission, with strategies aiming at reducing dephasing processes (such as quasi-resonant pumping or the application of a static electric field), will likely permit to generate highly-indistinguishable photons on demand.
From a technological point of view, our strategy for positioning single W centers is fully compatible with the geometry of photonic crystal cavities. In order to fully benefit from their large Purcell-enhancement factor, one has to implement means enabling a fine control of the spectral detuning between W’s ZPL and the high-Q cavity mode. In this context, electrical and strain tuning of the ZPL have been recently demonstrated for G centers. 48, 49

### v2

# Bright and pure single-photon source in a silicon chip by nanoscale positioning of a color center in a microcavity

@@ Abstract
We present an all-silicon source of near-infrared linearly-polarized single photons, fabricated by nanoscale positioning of a color center in a silicon-on-insulator microcavity. The color center consists of a single W center, created at a well-defined position by Si+ ion implantation through a 150 nm-diameter nanohole in a mask. A circular Bragg grating cavity resonant with the W’s zero-phonon line at 1217 nm is fabricated at the same location as the nanohole. By Purcell enhancement of zero-phonon emission, we obtain a photon count rate of Mcounts/s at saturation under above-gap continuous-wave excitation with a Debye-Waller factor of . A clean photon antibunching behavior is observed up to pump powers ensuring saturation of the W’s emission ( at ), evidencing that the density of additional parasitic fluorescent defects is very low. We also demonstrate the triggered emission of single photons with purity under weak pulsed laser excitation. At high pulsed laser power, we reveal a detrimental effect of repumping processes, that could be mitigated using selective pumping schemes in the future. These results represent a major step towards on-demand sources of indistinguishable near-infrared single photons within silicon photonics chips.
@@ Introduction
The generation, processing, and detection of single photons have enabled pioneering investigations in quantum physics, paving the way to groundbreaking technologies. Silicon quantum photonics, namely the processing of photonic quantum states in a silicon chip, is at the forefront of this revolution [42, 47]. However, the potential of silicon photonics for large-scale quantum applications is limited by the difficulty to deterministically generate arbitrary large entangled photon states on-chip. A key requirement to overcome this limitation is to embed sources of indistinguishable single photons operating on demand. While integrated sources of entangled photon pairs based on spontaneous four-wave mixing have enabled major demonstrations [47], these sources suffer from an inherent interplay between efficiency and multi-photon pairs emission. Alternative approaches rely on fiber-coupling of an external source [43, 5], and hybrid integration of quantum emitters [10, 14, 24], albeit with limited efficiency due to coupling losses.
@@
In this context, near-infrared color centers in silicon have emerged as a promising path towards deterministic single-photon sources (SPS) circumventing coupling losses. Due to their emission wavelength in the telecom bands, these point defects also appear as potential sources for chip-based quantum communications. Since the observation of photon antibunching for single color centers in 2020 [35, 19], tremendous efforts have been devoted to optimizing their elaboration processes [3, 20, 49, 2, 1], and exploring their photophysical properties [11, 4, 17, 33, 13]. In order to build a SPS, it is necessary to embed the emitter in a photonic cavity to achieve on-demand emission of single photons in a well-defined optical mode using the Purcell effect [32]. Several single-mode sources in silicon have been demonstrated, based on a T center [22, 23], G center [38, 25], G⋆ center [34, 31], and erbium dopant [16].However, in spite of remarkable improvement of emission properties owing to a high Purcell factor, these sources suffer from significant limitations. The T center and erbium dopant are inherently slow emitters [22, 23], the G center has a poor quantum efficiency [28, 12, 38], and the fabrication process of the G⋆ center is not selective [11], resulting in poor purity due to the presence of parasitic defects [34, 31].
@@
In recent work, Purcell enhancement has been demonstrated for an ensemble of W centers in circular Bragg grating (CBG) cavities in a silicon-on-insulator (SOI) film [29, 44]. On one hand, these results have positioned the W center as a promising candidate to build a SPS, owing to its high quantum efficiency () and relatively fast zero-phonon emission ( ns) in bulk silicon [29]. On the other hand, it confirms that CBG cavities, which have been widely used in other material systems [30, 8, 46, 26, 21, 15, 18], are also attractive for building SPS on SOI chips. However, due to a small mode volume, the position of the emitter in the cavity must be controlled accurately to ensure optimal exploitation of the Purcell effect and reproducible fabrication of the SPS.
@@
Here, we present a SPS fabricated by nanoscale positioning of a single W center in a SOI CBG cavity. An array of W centers is created by Si+ ion implantation through a mask patterned with nanoholes [20], then an array of CBG cavities centered at the same location as the nanoholes and resonant with the W zero-phonon line (ZPL) is fabricated. The optical properties of a cavity-coupled W center are studied by photoluminescence (PL) spectroscopy and photon correlation measurements under above-gap continuous-wave (CW) and pulsed excitation. The purity of the source is analyzed as a function of the laser excitation power.
@@
The W centers are created by a well-known combination of Si+ implantation and thermal annealing [48, 6]. Similarly to previous work [20], the implantation is performed through a mask patterned with nanoholes whose size is tuned to locally create single W centers. A 480 nm-thick mask of polymethyl methacrylate (PMMA) is deposited onto a commercial piece of SOI wafer, arrays of nanoholes defined by electron beam lithography are etched into the mask, and the sample is broad-beam implanted with Si+ ions. The PMMA mask is then removed by stripping with acetone and the sample is annealed at 200 °C for 30 minutes in N2 atmosphere to activate the emitters (Figure 1(b)), whose microscopic structure - consisting of three interstitial Si atoms [4] - is shown in the inset. For a nanohole diameter of 150 nm, single W centers are observed with a yield of 40% with low background emission () and average count rate of kcps. A characterization of the emitters in unpatterned SOI is presented in Section 3 of the Supporting Information (SI).
CBG cavities, consisting of a disk surrounded by four grating rings, are fabricated by electron beam lithography and reactive ion etching at the same location as the nanoholes (Figure 1(c)). After fabrication, the sample is inspected by scanning electron microscopy to confirm successful fabrication according to design parameters (see Figure S9 in the SI). A map of the electromagnetic field in the central disk of the cavity is shown in the inset. Note that this method differs from the more common strategy of determining the position of preselected emitters by PL mapping, that is more demanding to implement and where the positionning accuracy is limited by optical resolution [31].
@@
A sectional view of a CBG cavity containing a single W center is schematically represented in Figure 1(d), with the relevant dimensions. The red double-arrow represents the orientation of the transition dipole along the direction [4]. It is well-known for CBG cavities that the Purcell enhancement factor , defined as the ratio between the emission rate in the cavity and bulk semiconductor, is heavily dependent on the position of the emitter [41]. This sensivity was investigated for our emitter-cavity system using numerical finite-difference time-domain simulations (FDTD, see Section 4 in the SI). Figure 1(e) shows the Purcell enhancement factor for the ZPL as a function of the distance between the center of the cavity and the position of the dipole, in the coordinate system defined in Figure 1(d). A maximum enhancement factor of 13 is obtained at the center of the cavity. decreases rapidly against and reaches 0.3 for nm, emphasizing the requirement for high positioning accuracy. Considering the simulated lateral spread of the ions implanted through a nanohole of 150 nm diameter, we estimate that the W centers have near-unity probability to be located less than 120 nm away from the center of the cavity, i.e. in the vicinity of the main antinode of the electromagnetic field where Purcell enhancement is maximal (see Section 2 in the SI).
@@
The collection efficiency for a cavity-coupled W center was also investigated using FDTD, considering the configuration of our measurement setup. The fraction of dipole emission radiated through a numerical aperture of 0.65 at normal incidence is plotted against in Figure 1(e). A collection efficiency of 51 is obtained at the center of the cavity. Note that this value could be increased to nearly with the addition of a high-reflectivity mirror below the cavity [36]. The graph also reveals the robustness of the collection efficiency against spatial detuning (with for nm for each three coordinates), except for the "pit" around = 120 nm where the emitter is at a node of the electromagnetic field. This feature makes CBG cavities an attractive microstructure to easily increase the photon detection rate from quantum emitters in SOI. Similarly, the collection efficiency is also found robust against spectral detuning (see Figure S7 in the SI).
@@
After fabrication, the sample is characterized by mapping the PL intensity in the emission range of W centers (1200-1300 nm). In the following, we focus on cavities fabricated over nanohole arrays of 150 nm diameter, where single W centers have been identified after the annealing step. On PL maps, the cavities populated by W centers appear as bright spots on a dark background. The number of W centers formed by localized ion implantation being a random variable, finding cavities containing a single center can be a tedious task. Here, we use polarization-selective PL mapping to pre-select potential cavities. As the emission of the W center at normal incidence is linearly polarized along either the [110] or [10] crystal axes of silicon, corresponding to horizontal and vertical polarization in free space in our experimental configuration, single W centers appear as H- or V-polarized spots in PL maps. By recording H- and V-polarized PL on separate maps, potential cavities are identified by polarization contrast. Four pairs of cross-polarized PL maps are displayed in Figure 2(a). In each pair of maps, one or several bright spots exhibit a clear polarization contrast. By measuring the photon autocorrelation function in the Hanbury Brown and Twiss configuration, the presence of a single W center in a given spot is revealed by the observation of strong photon antibunching (<0.5). In practice, antibunching is systematically observed for spots with more than 90% contrast, which is the case for 6 of the 40 spots displayed in the figure (indicated by white ellipses). Thus, we estimate that the end-to-end efficiency of the fabrication process is of the order of .
In the following, we will focus on the spot labeled as W-1 in the map set I, whose polarization diagram (Figure 2(b)) evidences linear polarization with a contrast of .
The single-photon emission properties of the W-1 spot under above-gap CW excitation are displayed in Figure 3 (see Section 1 in the SI for details about the measurement configuration). A high-resolution PL map under strong pumping reveals a maximum intensity exceeding 1.2 Mcps when the PL emission is optimally collected (Figure 3(a)). The power characteristic of the emitter evidences a clear saturation behavior, with a count rate of Mcps at saturation. This value, which is 400 times larger than the average count rate obtained for single W centers in planar SOI, is remarkably high compared to previously published results for other Si color centers (See Table 1 for a comparison). We will show in the next paragraph that this increase results from Purcell enhancement of zero-photon emission and high collection efficiency. To confirm that the PL is emitted by a single W center, the photon autocorrelation function is measured. A clean antibunching dip is observed without any correction for detectors dark counts or background emission, with at (Figure 3(c, d)) and at (Figure S10 in the SI), evidencing high-purity single-photon emission up to pumping powers ensuring saturation of the emitter. These values are among the best reported values for Si color centers, as shown in Table 1, indicating that the source is nearly free of additional fluorescent defects. The autocorrelation histogram also exhibits a bunching behavior () at non-zero delay. This phenomenon, already reported for single W centers [4, 20], is attributed to a metastable state coupled to the excited state of the emitter. It is also worth mentionning that, as the measurements are performed under strong pumping, the width of the antibunching dip is defined by the average repumping time and does not correspond to the lifetime of the excited state [4].
| Em. | Cav. | (kcps) | (ns) | Pos. | Ref. | |
| T | 1D PhC | 20 | 169 | 0.07 | No | [22] |
@@
To understand cavity effects quantitatively, we investigated the optical properties of the W-1 center and the spectral resonance of the emitter-cavity system. The PL spectrum of the emitter over the 1140-1450 nm spectral range reveals that the ZPL accounts for of the PL intensity (Figure 4(a)). This value should be compared with 39% for isolated W in unpatterned SOI, which unveils selective enhancement of the ZPL by the cavity. To confirm this hypothesis, the modal properties of the cavity have been probed by performing reflection spectroscopy under white light illumination. The reflectivity curve features a dip around 1218 nm fitting to a Lorentzian profile with a full width at half maximum (FWHM) nm, corresponding to a quality factor (Figure 4(b)). This value is in good agreement with the quality factor estimate of 167 obtained from FDTD simulations. A comparison between the reflectivity curve and a high-resolution PL spectrum shows that the CBG cavity mode is well-matched to the ZPL of the W-1 center. We conclude that the selective enhancement of the ZPL is induced by (1) efficient coupling of zero-phonon emission into the cavity mode, whose narrow far-field radiation pattern ensures high collection efficiency, and (2) possibly acceleration of zero-phonon emission by Purcell effect.
@@
To estimate the magnitude of the Purcell effect, we study the recombination dynamics of the W-1 center. The PL decay of the emitter under pulsed excitation at 17.5 W with a 5 MHz repetition rate is displayed in Figure 4(c). The intensity decays nearly mono-exponentially with a characteristic time ns. This value should be compared with the characteristic time of 34 ns obtained for ensembles of W centers in unpatterned SOI [6, 29, 44]. The acceleration of the decay compared to this reference value has two possible origins: Purcell enhancement of zero-phonon emission, and/or activation of extra non-radiative recombination channels due to the interaction between the emitter and its environment. This second possibility cannot be neglected a priori as a significant lifetime dispersion has been reported for single W centers in 60 nm-thick SOI [4], although this dispersion can likely be attributed to surface effects in such a thin SOI layer. A detailed analysis of the photon budget taking into account the setup transmission provides a lower-bound estimate of 1/18 ns-1 for the zero-phonon emission rate (see Section 5 in the SI). By comparison with a previous estimate of 1/134 ns-1 in bulk Si [29], we conclude that the zero-phonon emission rate is increased by a factor . Therefore, the acceleration of the PL decay results predominantly (and possibly entirely) from the Purcell effect. Additionally, according to Figure 1(e), we can also conclude that the W-1 center is positioned less than 100 nm away from the cavity center.
@@
The bright emission of antibunched photons under CW excitation unveils the strong potential of W centers for building a deterministic SPS in SOI. Along this line, we have explored the triggered emission of single photons under pulsed excitation. Figure 5(a) shows uncorrected autocorrelation histograms obtained for different pump powers for a 5 MHz repetition rate. The normalized area as a function of time delay , where is the laser repetition period, of each peak is given by the coordinate of the squares on the right axis. In the low power regime, the peak at zero-delay has an area , demonstrating the triggered emission of single-photon pulses with purity. As for the measurement under CW excitation, a bunching behavior is observed (>1), indicating photoblinking of the source. Note that this effect is a rather common issue for solid-state quantum emitters and is expected for W, as this color center possesses a metastable excited state [4]. The area of the peaks follows an exponential law given by:
@@
where and are the average times of stay of the emitter in the "bright" and "dark" states, respectively [40]. For = 0.5, 4.4, and 17.5 W, = , , and ns, and = , , and ns, pointing towards an acceleration of the blinking with increasing power. Interestingly, the peaks for have a smaller area than expected from the model in the low-power regime. This effect, which has been reported for single InAs QD under above-gap excitation [39], may be related to a two-step transition from the ground state to the excited state, which is indeed predicted by ab-initio calculation [4].
@@
The area of the central peak increases with power, as shown in Figure 5(b). For = 17.5 W, the area of the peak reaches , meaning that about two photons are emitted on average per excitation pulse. A zoom on the peak with 20 ps time bins (Figure 5(c)) reveals an antibunching dip with a characteristic time of 60 ps (that disappears in Figure 5(b) due to rebinning). This antibunching behavior indicates that the two photons are emitted successively. An explanation is that, under strong above-gap pumping condition, a second exciton can be captured after the emission of a first photon, as reported for other quantum emitters [7]. This phenomenon is not surprising here, due to the optical injection of a large population of carriers by the pump pulse in the central disk of the cavity. As for QD-SPS, we expect to overcome this limitation through the implementation of resonant or quasi-resonant pumping schemes.
@@
In spite of promising results under CW excitation, the generation of single photon pulses remains relatively inefficient. At a repetition rate of 5 MHz, the PL count rate reaches 52 kcps at saturation, corresponding to a first-lens efficiency of a few percents (see Section 5 of the SI). The main factors limiting this value at this stage are easily identified: (1) the emitter is optically active only 17% of the time, which reduces its average efficiency by a factor of almost 6; (2) our CBG cavity funnels only half of the emission towards our collection optics; (3) non-radiative recombination limits the emitter’s quantum efficiency. For a single W center ideally placed at the center of a CBG cavity, the probability to generate a ZPL photon upon excitation is expected to reach 90% at most [29].
While point (1) and the issue of blinking has already been discussed, we see that points (2) and (3) can be radically improved by replacing our low-Fp CBG cavity by a high-Fp waveguide-coupled photonic crystal cavity, such as a nanobeam cavity [45, 22, 23]. As the mode volume is similar for PhC and CBG cavities, the nanoscale positioning method is expected to apply equally well for both types of cavity. It should also be noted that the accuracy of the positioning can be further improved by increasing the Si+ dose and decreasing the holes diameter. Proper engineering of light extraction towards the waveguide mode and a large but realistic Purcell factor of 150 would bring the zero-phonon emission rate above 109 s-1. This would ensure that the quantum efficiency of the ZPL, the spontaneous emission factor and the efficiency of the waveguide-coupled SPS are all near unity. Single photon coherence and indistinguishability are also required properties for applications to quantum simulations and computing. Until now, Hong-Ou-Mandel interference have only been demonstrated over relatively narrow time-windows, much shorter than the emitter lifetime [27]. Combining such a large Purcell acceleration of zero-phonon emission, with strategies aiming at reducing dephasing processes (such as quasi-resonant pumping or the application of a static electric field), will likely permit to generate highly-indistinguishable photons on demand.
From a technological point of view, our strategy is compatible with the integration of multiple single-photon sources connected to the same photonic circuit for quantum applications. In a photonic chip comprising all processing elements and single-photon detectors, W centers spots could be created at defined positions in a dedicated area of the chip. Single W centers would then be mapped and connected to the circuit by fabrication of waveguide-coupled PhC cavities. In order to fully benefit from their large Purcell-enhancement factor, one has to implement means enabling a fine control of the spectral detuning between W’s ZPL and the high-Q cavity mode. In this context, electrical and strain tuning of the ZPL have been recently demonstrated for G centers [9, 37].